# Predictive Control and Regret Analysis of Non-Stationary MDP with Look-ahead Information

### v1

@@ 3.2 Diameter
The diameter of MDP is commonly used in reinforcement learning [Gajane et al., 2018; Wu et al., 2022]. We extend the definition to non-stationary MDP.
@@
Given a non-stationary MDP, time $t$, state $s$, policy $\pi$, and a sequence of states $\{s^{(i)}\}_{i}$, define
$d_{t}(s,\pi,\{s^{(i)}\}_{i})=\inf\{\tau:\tau>0,s_{t+\tau}=s^{(t+\tau)}|s_{t}=s\},$
where $\{s_{t+\tau}\}_{\tau=0}^{\infty}$ is generated by policy $\pi$, and let $T(s^{\prime}|\pi,s)$ denote the minimal travel time from $s$ to $\{s^{(i)}\}_{i}$ starting at any $t$.
$T(\{s^{(i)}\}_{i}|\pi,s)=\sup_{t}\left\{\mathbb{E}[d_{t}(s,\pi,\{s^{(i)}\}_{i})]\right\}.$ (7)
Let $s_{i}^{*}:=\arg\max_{s}V_{i}^{*}(s)$. Define the diameter of a non-stationary MDP as
$D=\max_{s}\min_{\pi}\left\{T(\{s_{i}^{*}\}_{i}|\pi,s)\right\}.$ (8)
@@
For a stationary MDP, The diameter as defined in 4 is upper bounded by the conventional diameter definition for stationary MDP as in Wu et al. [2022].
The proof of Proposition 3.2 is straightforward, as the $\{s_{t+i}^{*}\}_{i}$ is constant for all $t,i$ in time-invariant MDP.
@@
The proof of Proposition 3.3 is deferred to Appendix B.
@@ 4.1 Algorithm Design
where the learner optimize on the forecast of reward and transitional probability, instead of the ground-truth as in (3). The learner picks action $a$ such that
@@
The algorithm is simple and intuitive. In line 3 of Algorithm 1, we forecast the system dynamics of the future $k$ steps. In line 4, we pick the first action that maximizes the reward in the future $k$ steps. The algorithm design is inspired by model predictive control [García et al., 1989], and we try to optimize the performance of the learner within the prediction horizon and decide the action by a dynamic programming style algorithm.
@@ 4.2 Regret Guarantee
In this section, we introduce our bound on regret as defined in (2).
@@
For any non-stationary MDP satisfying 1, Algorithm 1 achieves a regret of
$\begin{split}\mathcal{R}(\mathrm{MPDP})\leq&T\gamma^{\lfloor k/J\rfloor}D+2T\epsilon_{0}+4T\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)\\ &+2T\delta_{0}D+4T\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor J+j}D\right),\end{split}$
@@
In the error-free setting (the prediction does not have error), Theorem 4.1 simplifies to the following corollary.
@@
We observe that, in the error-free case (Corollary 4.2), the regret depends linearly on the time horizon $T$ and diameter $D$ and decays exponentially with the prediction horizon $k$ when 1 is satisfied. In particular, Algorithm 1 with a $\log$-prediction horizon $k=O(\log T)$ will obtain a regret sublinear in $T$. This means that predictions, even if a short horizon, are powerful in the sense that it leads to sublinear regret without any assumptions on the variation budget.
In the setting of inaccurate prediction (Theorem 4.1), we observe that regret grows linearly with the prediction error $\epsilon_{\ell},\delta_{\ell}$. It is important to note that the sensitivity to $\epsilon_{\ell},\delta_{\ell}$ decays exponentially in $\ell$, meaning that the regret is more sensitive to the prediction error of the near horizon than the long horizon. This means that even if the prediction error increases as $\ell$ increases, as long as such an increase is subexponential, using predictions in the far future with potentially large errors still has a positive impact on the overall peformance.
@@
We briefly outline the steps of the proof of Theorem 4.1 in Section 5. The full proof is deferred to Section E.
@@ 5 Proof Outline
We split the proof outline into three separate steps. In the first step, we show how much the Bellman operator increases the error of estimating value function by using the reward and transitional probability forecast, rather than the ground-truth system dynamics. In the second step, we bound the error of the estimated $Q$ function using the step-wise error bound of the Bellman operator. With the error of $Q$ function bounded, if the learner makes a mistake using the estimated $Q$ function, the maximal loss caused by that mistake can be bounded. Therefore, we can bound the maximal regret in the last step.
Step 1: One step error bound. As Algorithm 1 takes a greedy approach to optimize the reward within the prediction horizon of length $k$, we first need to approximate the optimal value function $V_{t}^{*}$ within the $k$ steps.
Let $\hat{P}_{t+\ell|t}$ denote the predicted transition probability for $(t+\ell)$th step at time step $t$, and let $\hat{r}_{t}$ denote the predicted reward function. Let $\hat{\psi}_{t}^{k}$ denote the vector of the maximal expected reward at time $t$ for the next $k$ steps,
$\hat{\psi}_{t}^{\ell}(s)=\begin{cases}0,&t+\ell>T\text{ or }\ell>k,\\ \max_{a}\Big{\{}\hat{r}_{t+\ell|t}(s,a)+\mathbb{E}_{s^{\prime}\sim\hat{P}_{t+\ell|t}(\cdot|s,a)}\hat{\psi}_{t}^{\ell+1}(s^{\prime})\Big{\}},&t+\ell\leq T,\ell\leq k.\end{cases}$ (11)
Similarly, let $\tilde{\psi}_{t}^{k}$ denote the vector of the maximal expected reward assuming that the forecast is completely accurate.
$\tilde{\psi}_{t}^{\ell}(s)=\begin{cases}0,&t+\ell>T,\\ \max_{a}\Big{\{}r_{t+\ell}(s,a)+\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,a)}\tilde{\psi}_{t}^{\ell+1}(s^{\prime})\Big{\}},&t+\ell\leq T,\ell\leq k.\end{cases}$ (12)
@@
Proof:   We proceed by proof of induction, if $k=0$, the equality is trivial. The induction step directly follows from (3) and (12).
@@
In order to bound the error between $\hat{\Psi}_{t}$ and $Q_{t}^{*}$, we need to bound two pairs of the difference: the difference between $\tilde{\Psi}_{t}$ and $Q_{t}^{*}$, and the difference between $\tilde{\Psi}_{t}$ and $\hat{\Psi}_{t}$. The details of those steps are deffered to Lemma D.3 and Lemma D.4 in Appendix D, respectively. With an error bound of our approximated $Q$ function, we can upper bound the loss of reward by each mistake made by the algorithm.
@@
Let $a$ be the action picked by Algorithm 1 with prediction error as defined in 1 at the $t$-th time step at state $s$, and let $a^{*}$ be the optimal action at time $t$, then,
$\begin{split}Q_{t}^{*}(s,a^{*})&-Q_{t}^{*}(s,a)\leq\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+1}^{*}\right)+2\epsilon_{0}+2\delta_{0}D+4\sum_{i=0}^{\lfloor k/J\rfloor-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)\\ &+4\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right).\end{split}$
The proof of Corollary 5.3 is deferred to Appendix D.
Step 3: Bounding regret Since the error of $Q$ function is bounded at every step, we can upper bound the regret $R$ by bounding the telescoping sum of $\mathbb{E}[\sum_{t}\left(Q_{t}^{*}(s_{t},a_{t}^{*})-Q_{t}(s_{t},a_{t})\right)]$. By bounding the error at each step, we obtain the error bound in Theorem 4.1.
@@ 6 Simulation
In this section, we show the result of two simulations: the queueing system and electric vehicle (EV) charging.
@@ 6.1 Queueing system
In the first experiment, we simulate a queueing system based on the setup provided in Example 1. Specifically, we consider a representative example of 3 servers whose service rates $\{\mu_{i}\}_{i=1,2,3}$ are 100, 10, 1, respectively, with varying load $\lambda_{t}$ fluctuating from 10 to 100.
@@
In our first experiment, we compare regrets of different lengths of the prediction horizon $k$ and the Fast Available Server (FAS) and ratio-of-servicerate-thresholds (RSRT) algorithm. FAS is a popular algorithm frequently used in practice, which sends any available job immediately to the fastest available server. RSRT is a threshold policy where a job is routed to the fastest among the available server only if the queue length exceeds a predetermined threshold [Larsen and Agrawala, 1983]. It has been proven to be the optimal policy in the two-server setting. We compute regret where the optimal policy has full knowledge of the transitional probability. For each $k\in\{1,\dots,15\}$, we run 200 trials and record the average regret for each $k$ value. The optimal policy we compute our regret from is the policy that is computed knowing all future transitional probability and reward functions. As shown in Figure 1(a), the arrival of jobs fluctuates periodically. The queue length of FAS is consistently the longest throughout the time horizon, and RSRT has a queue length similar to that of MPDP with $k=8$. MPDP with $k=12$ has the shortest queue length throughout most of the time steps. As shown in Figure 1(b), our algorithm outperforms both benchmarks with $k\geq 8$. Specifically, we see a decay in log-scale of regret. However, as we have $J$-stage contraction, the regret does not necessarily decrease with every increase in $k$. After $k=8$, the regret of MPDP consistently outperforms the benchmark algorithms.
(a) Average queue length for Algorithm 1 with different $k$ values. The shaded area is one standard deviation above and below the mean.
@@
(b) Regret (in log scale) of MPDP with different $k$ values. The red solid line shows the mean of the regret, and the shaded area shows the confidence interval .
Figure 1:
@@
In the second simulation, we fix $k=10$ and examine the relationship between the magnitude of the prediction error and regret. Although the learner can still forecast the system dynamics in the future, the predicted arrival rate of jobs $\hat{\lambda}_{t}:=\lambda_{t}+\mathcal{N}(0,\sigma_{t,\ell})$, where $\mathcal{N}(0,\sigma_{t,\ell})$ is a Gaussian noise, and $\sigma_{t,\ell}$ is the variance. In this simulation, we assume $\sigma_{t,\ell}$ to be constant of $t,\ell$. Therefore, the variance of the prediction error does not increase with the distance of the forecast into the future. As shown in Figure 2(a), the regret initially remained minimal and increased linearly after variance reaches 8, as shown in Theorem 4.1.
(a) Regret vs. the magnitude of error. As the variance of prediction error increases, both the average regret (solid line) and the variance of regret samples also increase.
@@
(b) Regret vs. the growth rate of variance of the prediction error with respect to the prediction horizon.
Figure 2:
@@
Most practical forecast are usually more accurate for closer future than distant future. In applications like wind power generation, it has been shown that the accuracy of forecasts decreases at a linear rate with respect to the distant in the future [Qu et al., 2013]. Therefore, in the third simulation, we fix $k=10$ and the initial variance of the prediction error to be $1$ and examine the relationship between regret and growth rate of variance with respect to the prediction horizon $g$. More specifically, for each forecast, we fix the variance $\sigma_{t,0}=1$, and $\sigma_{t,\ell}=g*\ell$. We are most interested in the case where $g>1$, as the variance tends to increase with respect to the distance to the future predicted by the forecast. As shown in Figure 2(b), the regret does increase with growth rate $g$, but the increase is relatively slow. Even for the case $g=4$, which indicates the variance increases by 4 times for every time step, thus reaching 40 times of the initial error at the end of the prediction horizon, the average regret merely increased 2 times. Indeed, by the expression in Theorem 4.1, the regret would never explode if the growth rate of variance is sub-exponential.

### v2

@@ 3.2 Diameter
The diameter of MDP is commonly used in reinforcement learning (Gajane18; wu22). We extend the definition to non-stationary MDP.
@@
Given a non-stationary MDP, time $t$, state $s$, policy $\bm{\pi}$, and family of sets $\{S^{(i)}\subset\mathcal{S}\}_{i}$, define
$d_{t}(s,\bm{\pi},\{S^{(i)}\}_{i})=\inf\{\tau:\tau>0,s_{t+\tau}\in S^{(t+\tau)}|s_{t}=s\},$
where $\{s_{t+\tau}\}_{\tau=0}^{\infty}$ is generated by policy $\bm{\pi}$, and let $\mathcal{T}(\{S^{(i)}\}_{i}|\bm{\pi},s)$ denote the minimal travel time from $s$ to the family of sets $\{S^{(i)}\}_{i}$ starting at any $t$, i.e.
$\mathcal{T}(\{S^{(i)}\}_{i}|\bm{\pi},s)=\sup_{t}\left\{\mathbb{E}[d_{t}(s,\bm{\pi},\{S^{(i)}\}_{i})]\right\}.$ (7)
Let $S_{i}^{*}:=\arg\max_{s}V_{i}^{*}(s)$ denote the set of states with the maximal value function at time step $i$. Define the diameter of a non-stationary MDP as
$D=\max_{s}\min_{\bm{\pi}}\left\{\mathcal{T}(\{S_{i}^{*}\}_{i}|\bm{\pi},s)\right\}.$ (8)
In this paper, the time horizon $T$ is finite. Therefore, when $i>T$, we have $S_{i}^{*}=\mathcal{S}$. As the agent will arrive at $\{S_{i}^{*}\}_{i}$ immediately after the $T$-th time step, we always have $D<T+1$ for any finite horizon MDP.
@@
For a stationary MDP, The diameter as defined in ˜3 is upper bounded by the conventional diameter definition for stationary MDP as in wu22.
The proof of Proposition˜3.2 is straightforward, as the $\{s_{t+i}^{*}\}_{i}$ is constant for all $t,i$ in time-invariant MDP.
@@
The proof of Proposition˜3.3 is deferred to Appendix˜B.
@@ 4.1 Algorithm Design
where the learner optimize on the forecast of reward and transitional probability, instead of the ground-truth as in equation 3. The learner picks action $a$ such that
@@
The algorithm is simple and intuitive. In line 3 of Algorithm 1, we forecast the system dynamics of the future $k$ steps. In line 4, we pick the first action that maximizes the reward in the future $k$ steps. The algorithm design is inspired by model predictive control (García89), and we try to optimize the performance of the learner within the prediction horizon and decide the action by a dynamic programming style algorithm.
@@ 4.2 Regret Guarantee
In this section, we introduce our bound on regret as defined in equation 2.
@@
For any non-stationary MDP satisfying ˜2, Algorithm 1 achieves a regret of
$\begin{split}\mathcal{R}(\mathrm{MPDP})\leq&T\gamma^{\lfloor k/J\rfloor}D+2T\epsilon_{0}+2T\delta_{0}D+2T\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)\\ &+2T\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor J+j}D\right),\end{split}$
@@
In the error-free setting (the prediction does not have error), Theorem˜4.1 simplifies to the following corollary.
@@
We observe that, in the error-free case (Corollary˜4.2), the regret depends linearly on the time horizon $T$ and diameter $D$ and decays exponentially with the prediction horizon $k$ when ˜2 is satisfied. In particular, Algorithm˜1 with a $\log$-prediction horizon $k=O(\log T)$ will obtain a regret sublinear in $T$. This means that predictions, even if a short horizon, are powerful in the sense that it leads to sublinear regret without any assumptions on the variation budget.
In the setting of inaccurate prediction (Theorem˜4.1), we observe that regret grows linearly with the prediction error $\epsilon_{\ell},\delta_{\ell}$. It is important to note that the sensitivity to $\epsilon_{\ell},\delta_{\ell}$ decays exponentially in $\ell$, meaning that the regret is more sensitive to the prediction error of the near horizon than the long horizon. Therefore, even if the prediction error increases as $\ell$ increases, as long as the increase is subexponential, using predictions in the far future with potentially large errors still has a positive impact on the overall performance.
@@
We briefly outline the steps of the proof of Theorem˜4.1 in Section˜5. The full proof is deferred to Appendix˜E.
@@ 5 Proof Outline
We split the proof outline into three separate steps. In the first step, we show that the estimated value function converges to the ground truth value function in the span semi-norm with increasing prediction horizon. In the second step, we bound the error of the estimated $Q$ function using the step-wise error bound of the Bellman operator. With the error of $Q$ function bounded, if the learner makes a mistake using the estimated $Q$ function, the maximal loss caused by that mistake can be bounded. Therefore, we can bound the maximal regret in the last step.
Step 1: One step error bound. As Algorithm˜1 takes a greedy approach to optimize the reward within the prediction horizon of length $k$, we first need to approximate the optimal value function $V_{t}^{*}$ within the $k$ steps.
Let $\hat{P}_{t+\ell|t}$ denote the predicted transition probability for $(t+\ell)$ -th step at time step $t$, and let $\hat{r}_{t}$ denote the predicted reward function. Let $\hat{\psi}_{t}^{k}$ denote the vector of the maximal expected reward at time $t$ for the next $k$ steps,
$\hat{\psi}_{t}^{\ell}(s)=\begin{cases}0,&t+\ell>T\text{ or }\ell>k,\\ \begin{split}&\max_{a}\Big\{\hat{r}_{t+\ell|t}(s,a)+\mathbb{E}_{s^{\prime}\sim\hat{P}_{t+\ell|t}(\cdot|s,a)}\hat{\psi}_{t}^{\ell+1}(s^{\prime})\Big\}\end{split},&t+\ell\leq T,\ell\leq k.\end{cases}$ (11)
Similarly, let $\tilde{\psi}_{t}^{k}$ denote the vector of the maximal expected reward with completely accurate forecasts.
$\tilde{\psi}_{t}^{\ell}(s)=\begin{cases}0,&t+\ell>T\text{ or }\ell>k,\\ \begin{split}&\max_{a}\Big\{r_{t+\ell}(s,a)+\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,a)}\tilde{\psi}_{t}^{\ell+1}(s^{\prime})\Big\}\end{split},&t+\ell\leq T,\ell\leq k.\end{cases}$ (12)
@@
Proof: We proceed by induction. If $k=0$, the equality is trivial. The induction step directly follows from equation 3 and equation 12.
@@
In order to bound the error between $\hat{\Psi}_{t}$ and $Q_{t}^{*}$, we need to bound two pairs of the difference: the difference between $\tilde{\Psi}_{t}$ and $Q_{t}^{*}$, and the difference between $\tilde{\Psi}_{t}$ and $\hat{\Psi}_{t}$. The details of those steps are deffered to Lemma˜D.3 and Lemma˜D.4 in Appendix˜D, respectively. With an error bound of our approximated $Q$ function, we can upper bound the loss of reward by each mistake made by the algorithm.
@@
Let $a$ be the action picked by Algorithm 1 with prediction error as defined in ˜1 at the $t$-th time step at state $s$, and let $a^{*}$ be the optimal action at time $t$, then,
$\begin{split}Q_{t}^{*}(s,a^{*})-Q_{t}^{*}(s,a)\leq&\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+1}^{*}\right)+2\epsilon_{0}+2\delta_{0}D+2\sum_{i=0}^{\lfloor k/J\rfloor-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)\\ &+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right).\end{split}$
The proof of Corollary˜5.3 is deferred to Appendix˜D.
Step 3: Bounding regret. Since the error of $Q$ function is bounded at every step, we can upper bound the regret $R$ by bounding the telescoping sum of $\mathbb{E}[\sum_{t}\left(Q_{t}^{*}(s_{t},a_{t}^{*})-Q_{t}^{*}(s_{t},a_{t})\right)]$. By bounding the error at each step, we obtain Theorem˜4.1.
@@ 6.1 Queueing system
In the first simulation, we simulate a queueing system based on the setup provided in Example˜1. Specifically, we consider a representative example of 3 servers whose service rates $\{\mu_{i}\}_{i=1,2,3}$ are 100, 10, 1, respectively, with time horizon $T=100$ and varying load $\lambda_{t}$ fluctuating from 10 to 100.
@@
In the first part of this simulation, we compare regrets of different lengths of the prediction horizon $k$ and the Fast Available Server (FAS) (Lin84), ratio-of-service-rate-thresholds (RSRT) algorithm (Ozkan14), and Non-Stationary Natural Actor-Critic (NS-NAC) (jali2025). FAS is a popular algorithm frequently used in practice, which sends any available job immediately to the fastest available server. RSRT is a threshold policy where a job is routed to the fastest among the available servers only if the queue length exceeds a predetermined threshold (Ozkan14). It has been proven to be the optimal policy in the two-server setting (Ozkan14). Notably, typical time-varying RL algorithm, such as NS-NAC, would not work in this example without a pretrained initial policy, as there are infinitely many states in the queuing problem, and the policy would keep exploring new states with longer queue lengths. For a proper comparison, we set the initial policy similar to RSRT with a small exploration probability for every action. We then compute regret where the optimal policy has full knowledge of the transitional probability. For each $k\in\{1,\dots,15\}$, we run 20 trials and record the average regret for each $k$ value. The agent has access to the predicted arrival rate of jobs with some Gaussian additive prediction error $\hat{\lambda}_{t}:=\lambda_{t}+\mathcal{N}(0,\sigma)$ with $\sigma\in\{0,1,2\}$. The optimal policy we compute our regret from is the policy that is computed knowing all future transitional probability and reward functions. The arrival of jobs fluctuates periodically with a $\sin$ function with magnitude 110 to simulate periodical change in demand. The queue length of FAS is consistently the longest throughout the time horizon, and RSRT has a queue length similar to that of MPDP with $k=8$. MPDP with $k=12$ has the shortest queue length throughout most of the time steps. As shown in the rest of Figure˜1, the proposed algorithm outperforms both benchmarks with $k\geq 8$ under noise-free setting and also outperforms both benchmarks with $k\geq 10$ with prediction error. Specifically, we see a decay in log-scale of regret. However, as we have $J$-stage contraction, the regret does not necessarily decrease monotonically with every increase in $k$.
(a) $\sigma=0$
(b) $\sigma=1$
(c) $\sigma=2$
@@
Figure 1: Figure˜1(a), Figure˜1(b), and Figure˜1(c) show the regret of MPDP under different additive prediction error $\mathcal{N}(0,\sigma)$. The red solid line shows the mean of the regret, and the shaded area shows the confidence interval.
@@
In the second part of the simulation, we fix $k=10$ and examine the relationship between the magnitude of the prediction error and regret more closely. Although the learner can still forecast the system dynamics in the future, the predicted arrival rate of jobs $\hat{\lambda}_{t}:=\lambda_{t}+\mathcal{N}(0,\sigma_{t,\ell})$. We first hold $\sigma_{t,\ell}$ to be constant of $t,\ell$ as in the first part. Therefore, the variance of the prediction error does not increase with the distance of the forecast into the future. As shown in Figure˜2(a), the regret initially remained minimal and increased linearly after variance reaches 8, as shown in Theorem˜4.1.
@@
Figure 2: Figure˜2(a) shows the average regret remains almost constant with variance of prediction error below 5 and starts to grow afterward. Figure˜2(b) shows the regret slowly increases with the growth rate of variance of the prediction error with respect to the prediction horizon.
@@
Most practical forecast are usually more accurate for closer future than distant future. In applications like wind power generation, it has been shown that the accuracy of forecasts decreases at a linear rate with respect to the distant in the future (Qu13). Therefore, in the third simulation, we fix $k=10$ and the initial variance of the prediction error to be $1$ and examine the relationship between regret and growth rate of variance with respect to the prediction horizon $g$. More specifically, for each forecast, we fix the variance $\sigma_{t,0}=1$, and $\sigma_{t,\ell}=g*\ell$. We are most interested in the case where $g>1$, as the variance tends to increase with respect to the distance to the future predicted by the forecast. As shown in Figure˜2(b), the regret does increase with growth rate $g$, but the increase is relatively slow. Even for the case $g=4$, which indicates the variance increases by 4 times for every time step, thus reaching 40 times of the initial error at the end of the prediction horizon, the average regret merely increased 2 times. Indeed, by the expression in Theorem˜4.1, the regret would never explode if the growth rate of variance is sub-exponential.
(a) $\sigma=0$
(b) $\sigma=1$
(c) $\sigma=2$
Figure 3: Figure˜3(a), Figure˜3(b), and Figure˜3(c) show the regret of MPDP under different additive prediction error $\mathcal{N}(0,\sigma)$ in a setting with a finite number of changes in MDP’s. The red solid line shows the mean of the regret, and the shaded area shows the confidence interval.
Lastly, we introduce a simplified version of the queueing problem in Example˜1. While keeping all other factors identical, we fix $\lambda_{t}\in\{30,130\}$ and switch $\lambda_{t}$ every 50 steps. By fixing $\lambda$ to only two potential values, we simplify the setting where a traditional RL algorithm with context detection might have an advantage. To incorporate more changes in MDP, we also extend the time horizon to $T=300$.
The result of the simulation is shown in Figure˜3. We introduce RL-CD proposed in Silva06 as a new benchmark. For a proper comparison, we also set the initial policy to be similar to RSRT, with a small exploration probability for every action. Due to the scarcity of changes, both RL-CD and NS-NAC perform worse than RSRT, which is provably optimal for queueing systems with constant $\lambda$. However, the proposed algorithm still outperforms RSRT with $k\geq 9$ in all three noise settings with $\sigma\in\{0,1,2\}$, demonstrating the effectiveness of MPDP when model prediction is available.